--++++ INAF queue observing, 2018 September 29-30
Observer: R. Carini, A. Gargiulo  
Telescope Operator:  D. G. Huerta 
Support Astronomer: O. Kuhn   on call 
Instrument: LUCI
  Summary 
At the sunset  the sky is  mostly clear , some clouds on the horizon  
instruments: MODS2 red grating is not working
 LBC-blue has some issues, LUCI1  works but have the same problems with the AGW, so ufficially LUCI1  works, but could be a lot of problems as the last days.  We do not use it.
Night duration (12\xB0 el twilight)  2:30- 12:15 UT = 
9 h 45m
Observing time: 
4h 50m  (50%)
Weather time loss: 
3h  55m (40%)
technical time loss 
1h m (10%)
  Data Summary 
	
		
			|  Project_RUN  | 
			 Objects  | 
			 Instrument  | 
			 Configuration  | 
			 ExpTime /Time_tot  | 
			 Images  | 
			 Notes  | 
		
	
	
		
			|  telluric for VIPD4  | 
			 HIP106817  | 
			 Luci2  | 
			 ls, zjspec  | 
			    | 
			    | 
		
		
			|  Cassar\xE1  | 
			 VIPD4  | 
			 Luci2  | 
			 ls, zjSpec  | 
			 2.0/2.0  | 
			 luci2.0059-0095  | 
			 completed  | 
		
		
			|  Polletta telluric  | 
			 HIP106817  | 
			 mos, zjSpec  | 
			 luci2.0099-0104  | 
			 good for the first slit  | 
		
		
			|  Polletta  | 
			 PLCKG073  | 
			 mos,zjPspec  | 
			 0.47h  | 
			 luci2.0109-0122  | 
			    | 
		
		
			|  Polletta telluric  | 
			 HIP13917  | 
			 zjSpec  | 
			    | 
			 luci2.0127-0129  | 
			 good for the second slit  | 
		
		
			|  Perna telluric  | 
			 HIP16111  | 
			 Hkspec  | 
			 luci2.0133-0136  | 
			    | 
		
	
  Details 
02:03 start with the  telluric for the target of  Cassar\xE1 program: VIPD4 
-  acq 
luci2.20180930.0046-0048
  -  sci 
luci2.20180930.0049.0049-0052
  -  seeing 0.96''on the dimm
 
 
02:23   target VIPD4 
-  acq 
luci2.0053-
  -  02:30 error message during the alignment
  -  02:34 try again, now it works
  -  Aladin gives problem in the readout the images,
  -  02:59 sci 
luci2.0059-0077
  -  03:25 seeing about 1.0''
  -  03:40 some bursts of seeing to 1.2''
  -  03: 50: seeing about 1.0''
  -  04:10 load the script for the second time 
luci2.0078-0095
  -  seeing about 0.9''-1.0''
  -  04:13 some bursts up to 1.2'' but the mean seeing is 1.0''
  -  04:48 seeing about 0.9''
  -  05:13 seeing 0.8''-0.9''
 
 
05:16 slew to Polletta program, telluric HIP106817 
-  05:18 preset failed
  -  05:24 send the script again
  -  acq 
luci2.0096-0098
  -  sci 
luci2.0099-0104
  -  seeing 0.9''-1.0''
 
 
05:41 slew to Polletta target 
-  05:43 checkin pointing
  -  05:45 sent again the script
  -  check pointing
  -  05:47 sent again the script
  -  acq 
luci2.0105-0108
  -  06:00 sci==0109-0122==
  -  seeing about 0.9''
  -  06:30 the seeing has a burst at 1.4'' and come back to the value 1.0''
 
 
06:35 TO closes the dome due to clouds
10:03 we open again
10:17 go to the telluric Antoniucci HD 216532
   *10:20 MOS error
11:00 we slew to telluric Polletta HIP 13917 
-  we take the images only for the second slit
  -  acq 
luci2.0124-0.126
  -  sci 
luci2.0127-0129
  -  seeing 0.9
 
 
11:10 slew to telluric Perna program 
 
11:30 slew to telluric again 
-  acq 
luci0130-0132
  -  sci 
luci.0133-0136
 
 
11::45 slew to target 
 
11:52 TO close the dome due to clouds
calibrations
Cassar\xE1 
-  flat 
luci2.0145-0149
  -  arc HKSpec 
luci2.0150-0153
 
 
-- %USERSIG{RobertaCarini - 2018-09-30}%